How many atoms of hydrogen are represented in 2 H2C6H8O4?
1
Expert's answer
2020-11-04T14:16:01-0500

"2H_2C_6H_8O_4"


Hydrogen atoms==>

"2(H_2 + H_8)\\
=2(2+8)\\
=2(10)
=20"


"\therefore" There are 20 Hydrogen atoms in the compound